Nana can paint a room in 6 hours,Harith can do the job in 4 hours,but Angela need 2 hours to paint the same room.How long does it take them to paint the room if they work together?
Answer by greenestamps(13198)



A standard algebraic solution....

Nana paints 1/6 of the room in 1 hour; Harith paints 1/4 of it in 1 hour; Angela paints 1/2 of it in 1 hour. So together in 1 hour the fraction of the room they paint is

1/6+1/4+1/2 = 2/12+3/12+6/12 = 11/12

So the number of hours it takes the three of them together to paint the room is 1/(11/12) = 12/11.

ANSWER: 12/11 hours

An alternative solution avoiding fractions....

The least common multiple of the three times is 12 hours.

In 12 hours, Nana could paint 12/6=2 of the rooms
In 12 hours, Harith could paint 12/4=3 of the rooms
In 12 hours, Angela could paint 12/2=6 of the rooms

So in 12 hours the three of them together could paint 2+3+6=11 of the rooms

So the number of hours it would take them to paint the one room is 12/11.